### v1

# SAM-LAD: Segment Anything Model Meets Zero-Shot Logic Anomaly Detection

Abstract
Visual anomaly detection is vital in real-world applications, such as industrial defect detection and medical diagnosis. However, most existing methods focus on local structural anomalies and fail to detect higher-level functional anomalies under logical conditions. Although recent studies have explored logical anomaly detection, they can only address simple anomalies like missing or addition and show poor generalizability due to being heavily data-driven. To fill this gap, we propose SAM-LAD, a zero-shot, plug-and-play framework for logical anomaly detection in any scene. First, we obtain a query image’s feature map using a pre-trained backbone. Simultaneously, we retrieve the reference images and their corresponding feature maps via the nearest neighbor search of the query image. Then, we introduce the Segment Anything Model (SAM) to obtain object masks of the query and reference images. Each object mask is multiplied with the entire image’s feature map to obtain object feature maps. Next, an Object Matching Model (OMM) is proposed to match objects in the query and reference images. To facilitate object matching, we further propose a Dynamic Channel Graph Attention (DCGA) module, treating each object as a keypoint and converting its feature maps into feature vectors. Finally, based on the object matching relations, an Anomaly Measurement Model (AMM) is proposed to detect objects with logical anomalies. Structural anomalies in the objects can also be detected. We validate our proposed SAM-LAD using various benchmarks, including industrial datasets (MVTec Loco AD, MVTec AD), and the logical dataset (DigitAnatomy). Extensive experimental results demonstrate that SAM-LAD outperforms existing SoTA methods, particularly in detecting logical anomalies.
Index Terms:
Anomaly detection, Anomaly localization, Zero-shot, Segment Anything Model, Keypoint Matching.
I Introduction
Figure 1: (a) Example of the structural anomalies. (b) The anomaly-free images of the category breakfast box and screw bag. (c) The anomaly score map of the Patchcore, GLCF, and SAM-LAD for logical anomaly detection.
In recent years, image anomaly detection techniques have been widely applied in industrial quality detection[1][2][3][4], anomaly segmentation[5], and medical diagnosis scenarios[6], aiming to detect abnormal data that are different from normal data within a sample image[7]. Since abnormal prior information is scarce, this task is commonly conducted within the framework of an unsupervised learning paradigm. Consequently, there has been increasing interest among scholars in researching unsupervised anomaly detection, and near-perfect results have been achieved, as evidenced by methods such as Pull&Push[8], PMB-AE[9], and Patchcore[10]. However, due to the limits of those scenarios, most anomaly detection methods currently focus on structural anomaly detection only can deal with even one object in a single scene at at a time, as illustrated in Fig. 1(a).
In many usual scenarios such as autonomous driving and surveillance systems, understanding the semantic context of the entire scene and detecting anomalies is essential. Therefore, Bergmann et al.[11] have proposed logical anomalies, which represent more complex abnormalities in the logical relationships between objects within the entire scene. Specifically, within a scene, anomalies that appear normal at the local level but violate geometric constraints or logical principles when considering global semantics are referred to as global logical anomalies. For example, in Fig. 1(b), anomaly-free images of the category breakfast box always contain exactly two tangerines and one nectarine that are always located on the left-hand side of the box. Furthermore, the ratio and relative position of the cereals and the mix of banana chips and almonds on the right-hand side are fixed. An anomaly-free screw bag contains exactly two washers, two nuts, one longscrew, and one short screw. Nevertheless, their logical anomalies could involve a missing object, an unexpected object, or incorrect ordering and combination.
For the more challenging global logical anomalies, most existing methods tailored for structural anomalies demonstrate catastrophic results, for instance, as illustrated in Fig. 1(c) with Patchcore. This is because structural or textural anomalies belong to lower-level anomaly types, which do not necessitate an understanding of the overall semantics of objects in the scene and only require local knowledge for anomaly detection. However, for higher-level global logical anomalies, relying solely on local perception to ascertain the normalcy of overall semantics is insufficient. Consequently, the performance of existing methods is significantly constrained.
Our previous work[12] proposed a reconstruction-based method for detecting multiple object anomalies to address logical anomalies. In [12], objects had fixed spatial relationships. We used semi-supervised learning on positive samples to regress object positions in the test samples and reconstruct their normal features. However, in more complex scenarios like [11], where objects lack fixed spatial relationships, exploring new methods is essential to address logical anomalies.
Currently, Yao et al. proposed GLCF[13], Guo et al. proposed THFR[14], and Zhang et al. proposed DSKD[15]. These state-of-the-art methods, based on reconstruction and knowledge distillation, achieve decent performance in detecting logical anomalies. For example, GLCF attempted to focus on anomalies in nectarines on the left-top side, as shown in Fig. 1(c). Unfortunately, GLCF exhibited disappointing results when encountering the screw bag. This is because, in the screw bag scenario, anomaly-free samples involve randomly placed objects, which requires a high level of contextual understanding and the exclusion of interference from diverse positive sample features during inference. Additionally, these methods are heavily data-driven, making them poorly generalizable and costly to retrain for new scenarios.
To address the above challenges: 1) multiple key objects in complex scenes. 2) variety of positive sample features. 3) poor generalization because of dependence on data-driven. We propose a novel framework called SAM-LAD, which can be plug-and-play in any scenario without training and even outperforms existing data-driven logical anomaly detection methods. Specifically, we ingeniously leverage the robust object segmentation capability of the Segment Anything Model (SAM)[16] to obtain the positional information of all key objects in the scene by setting segmentation thresholds. Then, we utilize a pre-trained DINOv2[17] backbone network as a feature extractor to extract features and employ nearest neighbor search to find the $k$ most similar normal samples to the query image. Following this, we implement the FeatUp[18] operation to upsample the feature maps and recover their lost spatial information. Combining the upsampled feature maps with the positional information obtained from SAM yields separate feature maps for each object. Subsequently, we propose a Dynamic Channel Graph Attention (DCGA) mechanism to effectively compress each object’s feature map into a single feature vector. Consequently, leveraging the proposed Object Matching Model (OMM), we match the feature vectors of the reference images with those of each feature vector in the query image. Finally, based on the matching results, we propose an Anomaly Measurement Model (AMM), which estimates the feature distribution of individual objects in the query image and the corresponding $k$ matched objects from the $k$ normal samples, thereby calculating the final anomaly score map. Noteworthy, thanks to the introduction of AMM, our framework not only excels in detecting global logical anomalies but also fulfills the requirements for detecting structural anomalies. Our framework is motivated by the following: when humans discern anomalies among multiple objects within intricate scenes, the most simplest and efficacious approach involves juxtaposing several normal samples with each object in the test sample individually. Through this meticulous comparison, inference regarding anomalous regions can be inferred without necessitating the establishment of a costly global semantic contextual comprehension within the model. The essence lies in identifying and aligning pivotal objects, thus accomplishing the entirety of the task. We evaluate the proposed framework on multiple commonly used benchmarks, and the experimental results demonstrate that our SAM-LAD achieves state-of-the-art performance. The main contributions of this paper can be summarized as follows:
  1. 1.
    We novelty introduce SAM into the field of logical anomaly detection, leveraging its strong segmentation prowess to obtain the positional information of each crucial object within complex scenes without training. With this positional information as the foundation, subsequent matching between objects can be facilitated.
  2. 2.
    We propose the DCGA mechanism, which, drawing from the principles of graph neural networks, can effectively compress the upsampled feature maps of individual objects while considering the logical dependencies among them, thereby preserving their most crucial features in the form of feature vectors. DCGA ensures the accuracy and efficiency of subsequent object matching.
  3. 3.
    We innovatively propose the OMM, which leverages the principles of optimal transport and the feature vectors obtained from DCGA to generate a matching score matrix between each pair of objects. Subsequently, utilizing the Sinkhorn algorithm to solve the matching score matrix yields the matching relationships between objects. Based on these relationships, anomalies within objects can be accurately identified.
  4. 4.
    We propose the AMM, which computes the feature distribution differences between a single object in the query image and the matched objects from positive samples. This model enables not only the detection of logical anomalies but also the identification of structural anomalies present within them.
II Related Work
II-A Semantic Segmentation
The field of semantic segmentation has witnessed significant advancements, with novel methods being developed that target improved precision and versatility. The Segment Anything Model (SAM)[16], a newly introduced framework, is revolutionizing object segmentation by leveraging both large-scale training datasets and innovative model architectures. Existing studies have explored various techniques, such as transformers with self-attention mechanisms[19], as well as methods that integrate diverse image contexts to enhance segmentation performance. Prior to SAM, notable segmentation models like Mask R-CNN[20] and DeepLab[21] have paved the way in extending segmentation capabilities through region proposals, feature pyramid networks, and atrous convolutions. SAM continues this trend of innovation by adopting a prompt-driven architecture that not only enhances traditional segmentation tasks but also explores uncharted territories in segmentation capabilities. Furthermore, interactive segmentation methods have previously been explored to address the ambiguity of object boundaries by enabling user inputs to guide segmentation predictions. However, SAM’s approach differs by providing a flexible way to specify segmentation prompts (e.g., points, boxes, masks) and generalize the segmentation to diverse scenarios.
In this work, we use SAM’s robust segmentation capabilities and generalization performance. By setting segmentation thresholds, we can effectively isolate the key objects to be detected from the scene. Consequently, our framework achieves a groundbreaking analysis of the entire scene, all based on zero-shot learning.
II-B Keypoint Matching
Keypoint matching is a fundamental task in computer vision, crucial for applications such as object recognition, 3D reconstruction, and image stitching. The long-standing interest in this problem has resulted in several approaches over the years. Early classical methods like SIFT[22] and ORB[23] laid the groundwork by detecting handcrafted keypoint descriptors. These algorithms focused on creating robust descriptors that were invariant to scale, rotation, and partial occlusion, which made them effective for a wide range of tasks, despite being sensitive to image illumination changes. In recent years, deep learning techniques have revolutionized keypoint matching by learning discriminative feature representations from large datasets. CNNs have been widely adopted in keypoint detection and matching, such as the SuperPoint model[24], which directly learns keypoint detection and descriptor generation in an end-to-end fashion. Deep learning-based approaches offer much-improved robustness and generalization compared to handcrafted methods. More recent advances employ attention mechanisms to improve keypoint matching under challenging environments. Networks like SuperGlue[25] harness graph neural networks along with self- and cross-attention mechanisms to establish robust keypoint correspondences. The introduction of transformers into keypoint matching has shown further potential in improving the model’s context understanding, enabling them to handle complex matching scenarios more proficiently.
Inspired by SuperPoint and Superglue, we consider each object in the scene as a keypoint, and compress the keypoint features into a single feature vector using the proposed dynamic channel graph attention (DCGA) mechanism. Subsequently, we compute the object-to-object matching relationship using the concept of optimal transport.
II-C Unsupervised Anomaly Detection
The core part of unsupervised anomaly detection algorithms lies in extracting prior knowledge of anomaly-free objects, serving as a criterion for anomaly detection tasks. Typically, current unsupervised anomaly detection approaches can be classified into two categories based on the type of anomalies they address: methods targeting structural anomalies, and methods targeting logical anomalies. Presently, the majority of anomaly detection methods are designed for structural anomalies, primarily using a robust feature extraction network to obtain high-level semantic features of a test image. The distance between features of test images and anomaly-free images is calculated, and the areas with large distances are determined to be anomalies. The existing methods usually use deep CNN models, such as ResNet[26] and EfficientNet[27], for feature extraction. For instance, Roth et al.[10] proposed Patchcore to detect anomalies on objects with a concise background. ResNet-50[26] pretrained on ImageNet[28] was selected as the feature extractor. However, these methods falter when detecting logical anomalies, as the local structures within the image appear normal. Hence, an increasing number of researchers are directing their focus toward logical anomaly detection. For instance, Zhang et al. have proposed DSKD[15], and Batzner has introduced EfficientAD[29]. Both methods are based on the Student-Teacher network and anomalies are obtained by comparing teacher and student differences. Leveraging a local-global branching approach, Yao et al. presented GLCF[13], while Bergmann et al. proposed GCAD[11] in a bid to understand the logical semantic constraints governing the entire image scene. Guo et al. presented THFR[14], which is based on template-guided reconstruction. These methods have exhibited promising efficacy in detecting certain logical anomalies, such as missing components and unexpected excess. However, their detection performance significantly drops when confronted with more complex logical anomalies, such as misordering, mismatches, and haphazard arrangement of objects.
The proposed SAM-LAD framework builds a zero-shot method without strenuous efforts to comprehend the logical semantics of the entire scene. By introducing SAM and proposing object matching model and anomaly measurement model, a global logical anomaly detection system was realized, perfectly addressing the limitations encountered by the existing methods mentioned above.
III Methodology
Figure 2: Overview of the proposed framework, which consists of 4 stages. The first stage is an offline operation, which is building an anomaly-free template features bank. The second stage is to extract a feature map from a query image. The next stage is to utilize SAM to further obtain object feature maps. The last stage involves matching the objects in the query image with those in the reference images one by one. Using the obtained matching relationships, we calculate anomalies within each object and obtain the final anomaly score maps.
III-A Architecture Overview
The data flow of the proposed SAM-LAD is depicted in Fig. 2, which consists of four stages. 1) On-boarding stage 1 is an offline operation. For all normal images $\mathbb{R}^{3\times H\times W}$, a pre-trained backbone as a feature extractor is first deployed to extract the features from all normal images. Subsequently, the extracted features are compiled and stored within a template features bank. 2) Stage 2 involves extracting features from a query image $I_{q}\in\mathbb{R}^{3\times H\times W}$ using the same feature extractor as stage 1. 3) Stage 3, for the feature maps $f_{q}$ extracted from a query image $I_{q}$, we retrieve its $k$-nearest normal feature maps $(f_{r}^{i})_{i\in{[1,k]}}$ in the template bank and their corresponding images $(I_{r}^{i})_{i\in{[1,k]}}$. We designate the $k$ normal images procured as reference images and in parallel with the query image, input $(I_{q},I_{r}^{i})_{i\in{[1,k]}}$ into the SAM to obtain the individual object mask. Concurrently, the feature maps $(f_{q},f_{r}^{i})_{i\in{[1,k]}}$ are subjected to a FeatUp operation, upsampling to 8$\times$. By combining the individual object masks with the upsampled feature maps, we obtain the object feature maps for both the query and reference images, respectively. 4) In Stage 4, the DGCA module is employed to encode and compress the object feature maps into object descriptor vectors. Subsequently, utilizing the proposed Object Matching Module, the two sets of object descriptor vectors are matched. Based on the matching results, the Anomaly Measurement Module computed the ultimate anomaly score map.
III-B Feature Extraction and Template Features Bank
The first and second stage of the proposed SAM-LAD is the extraction of feature maps. The same feature maps are later used for FeatUp operation. There are multiple options for extracting features. Recently, Vision Transformer (ViT) has exhibited remarkable performance in anomaly detection-related tasks[30] [31] due to its self-attention mechanism, enabling the model to attend to global information of the entire image when processing image patches. Therefore, we have adopted the pre-trained ViT-type DINOv2-S backbone[17], as the feature extractor for the proposed SAM-LAD. For a given image $I_{x}^{3\times H\times W}$, we denote the extracted feature maps $f_{x}^{C\times H^{{}^{\prime}}\times W^{{}^{\prime}}}$:
$f_{x}=F(I_{x}),$ (1)
where $F(\cdot)$ is DINOv2 feature extractor.
At initialization, we execute offline operations to construct a template features bank =$\mathcal{B}=\left\{B_{1},\cdots,B_{i},\cdots,B_{Z},\right\}$ using $F$ to extract $Z$ normal images in the all normal set $\mathbb{R}^{3\times H\times W}$, where $B_{i}$ represents the template feature map of the $i$-th normal sample. At inference, only the feature maps of the query image are extracted and we could reduce the template features bank using coreset subsampling method[32][33] to reduce inference time and memory usage.
III-C Image-level Nearest Neighbor Search
Given a query image $I_{q}\in\mathbb{R}^{3\times H\times W}$, we take the extracted feature map $f_{q}$ as the query to retrieve its correlated template. ImageNNS obtains the template with index $t$ by randomly selecting from the template candidates that are $k$ most similar templates to increase the robustness during the inference process. The template selection process could be formulated as follows:
$t=\operatorname{random}\left(\underset{\mathcal{S}\subset\{1,\cdots,Z\},|\mathcal{S}|=k}{\arg\min}\sum_{i\in\mathcal{S}}d\left(f_{q},B_{i}\right)\right),$ (2)
where $d(\cdot)$ denotes the images-level distance between input query feature map $f_{q}$ and template feature key $B_{i}$ by flattening them to vectors to compute Euclidean metric. $S$ is a subset of $\left\{1,\cdots,Z\right\}$ denotes the indexes of $k$ template candidates, which are the top-k nearest of the input feature map $f_{q}$. Compared with the traditional point-by-point search strategy, imageNNS not only ensures that the reference is completely normal but also improves search efficiency.
Based on the obtained index $t$, we acquire the $k$ closest reference images $(I_{r}^{1},I_{r}^{2},\cdots,I_{r}^{k})^{k\times 3\times H\times W}$ along with their corresponding feature maps $(f_{r}^{1},f_{r}^{2},\cdots,f_{r}^{k})^{k\times C\times H{{}^{\prime}}\times W^{{}^{\prime}}}$.
III-D Object Feature Map Generator
III-D1 FeatUp Operation
For a given query image $I_{q}$, we now possess $k$ pairs of feature maps $(f_{q},f_{r}^{i})_{i\in[1,k]}$. These deep feature maps already capture the semantics of the images. However, these feature maps lack spatial resolution, making them unsuitable for directly performing subsequent dense anomaly detection and segmentation tasks, as the model would aggressively pool information over large areas. Therefore, inspired by FeatUp[18], we have employed its pre-trained feed-forward JBU upsampler[18] to restore the spatial information lost in the existing deep feature maps while still retaining the original semantics. Considering the balance of computational resources, inference speed, and ultimate detection accuracy, we opt to upsample the original feature maps by 8$\times$ (further analysis in Section IV-F). Subsequently, we upsample each pair of feature maps to obtain the upsampled feature maps, which size is $(C\times 8H^{{}^{\prime}}\times 8W^{{}^{\prime}})$:
$((f_{q})^{{}^{\prime}},(f_{r}^{i})^{{}^{\prime}})=JBU(f_{q},f_{r}^{i}),i\in[1,k]$ (3)
III-D2 Object mask
We use SAM to generate object masks for the input image pairs $(I_{q},I_{r}^{i})_{i\in[1,k]}$. However, in actual applications, due to SAM’s superior segmentation performance, it often segments both overly detailed objects and those not intended for detection. For example, in Fig. 3(a), in the breakfast box category, a target object might be a mix of banana chips and almonds, but SAM segments individual banana chips and almonds(Fig. 3(b)), which is not what we intended. Thus, after obtaining the segmentation masks from SAM, it is necessary to further filter them to acquire the desired key object masks.
Figure 3: (a) A target object of breakfast box category, (b) Unexpected overly detailed objects mask.
Specifically, upon processing the image data through the SAM, we can obtain the area of each segmented object. By setting thresholds for the minimum and maximum area, we can effectively filter the segmented objects, thus isolating the key objects of interest. For each category of the scene, SAM filters are set to determine the minimum and maximum area thresholds, resulting in object masks. The key object of interest as illustrated in Fig. 4.
Figure 4: The anomaly-free image’s object mask of each category dataset from SAM.
After segmentation by SAM and filtration, we obtain the $k$ pairs of segmented object masks for the image pairs $(I_{q},I_{r}^{i})_{i\in[1,k]}$. For the segmented object mask from anomaly-free reference image $I_{r}$, the number of object classes within its object mask is invariably constant. For instance the category of the breakfast box, an anomaly-free image always contains two tangerines, one peach, a portion of cereals, and a mix of banana chips and almonds, in addition to the main breakfast box itself, totaling six objects. Another category of screw bag contains exactly two washers, two nuts, one longscrew, two screw heads, and one short screw, totaling eight objects. Hence, for the segmentation of object masks from the anomaly-free reference image $I_{r}$, we partition it into $N$ individual object masks. (For detailed $N$ values for each category, refer to the red numerical annotations of object masks in Fig. 4). For the segmented object masks from $I_{q}$, due to the presence of missing or additional objects, the number of object classes M in the segmented object masks varies. We partition it into M object masks. Ultimately, for the image pairs $(I_{q},I_{r}^{i})_{i\in[1,k]}$, we obtain $k$ pairs of M and N object masks, respectively.
III-D3 Object Feature Map
For the $M$ object masks from the query image $I_{q}$, we first use bilinear interpolation to scale them to a size of $(8H^{{}^{\prime}},8W^{{}^{\prime}})$. Then, we perform element-wise multiplication with the corresponding upsampled feature map $(f_{q})^{{}^{\prime}}$, yielding $M$ object feature maps $(M,C,8H^{{}^{\prime}},8W^{{}^{\prime}})$. Similarly, for the $N$ object masks from the reference image $I_{r}$, we perform the same operation to obtain $N$ object feature maps $(N,C,8H^{{}^{\prime}},8W^{{}^{\prime}})$. Ultimately, we have obtained $k$ pairs of object feature maps $(f_{q}^{obj},(f_{r}^{obj})^{i})_{i\in[1,k]}$.
III-E Object Matching
III-E1 Dynamic Channel Graph Attention
To facilitate the matching of objects between the query and reference images, we consider each object to be a keypoint and extract their respective feature map into feature descriptor vector. However, during the process of feature compression, it is inevitable that positional information will be lost. Additionally, the recalibration of channel weights, by enhancing or suppressing semantic information, renders the extraction of global information challenging. Therefore, the establishment of a Dynamic Channel Graph Attention (DCGA) module is proposed to enhance the responsiveness between objects and channels, which can explicitly capture the spatial dependencies of various objects to augment global representation. Specifically, the feature maps of all objects can be conceptualized as a graph structure. Within this concept, each individual object feature can be regarded as a vertex in the graph, and the relationships between these objects are seen as the edges. Moreover, within the feature map of each object, each individual channel can also be viewed as a vertex, with the interactions between these channels representing edges. A schematic of the DCGA mechanism is illustrated in Fig. 5.
Figure 5: Architecture of the DCGA module.
For a query image’s objects feature map $(M,C,8H^{{}^{\prime}},8W^{{}^{\prime}})$, the dimension of each object feature map is first squeezed to $C\times 1\times 1$ by a global max pooling (GMP) operation. Then, the object features in DCGA is a tensor of shape $M\times C$. Subsequently, a graph structure is employed to generate the weights for each vector.
This graph structure consists of two parts. The first part targets each object’s feature vector $(1,C)$. Specifically, two independent $C\times C$ matrices, that is, $A_{0}$ and $A_{1}$, constitute the adjacency matrices, representing the dependency relationships among channel vertices. $A_{0}$ is generated by a 1-D convolution layer and a Softmax function. It is a self-attention-based diagonal matrix designed to suppress irrelevant features. $A_{1}$ is a predefined identity matrix, representing only the vertex itself, and requires normalization. Thus, the adjacency matrix can be represented as:
$A=A_{0}\times A_{1}$ (4)
The second part pertains to all the object feature vectors $(M,C)$, where cosine similarity is utilized to calculate the similarity between each object feature, thereby deriving the object feature adjacency matrix $A_{2}$:
$A_{2}$ $=\begin{bmatrix}0&S_{12}&\cdots&S_{1j}\\ S_{21}&0&\cdots&S_{2j}\\ \vdots&\vdots&\ddots&\vdots\\ S_{i1}&S_{i2}&\cdots\ &S_{ij}\\ \end{bmatrix},$ (5)
$S_{ij}$ $=\frac{S_{ij}^{\prime}}{\sum_{k=1}^{M}S_{ik}},$
$S_{ij}^{\prime}$ $=\hat{v}_{i}\cdot\hat{v}_{j}=\sum_{d=1}^{C}\hat{v}_{id}\cdot\hat{v}_{jd}.$
Where $S_{ij}^{\prime}$ represents the similarity between two object vertex in the M and $S_{ij}$ is its normalization. Note that, when generating the adjacency matrix, the weights updating is based solely on the similarity between one object vector and others, self-similarity is set to zero, that is, the diagonal of $A_{2}$ is zero.
In a nutshell, the DCGA can be formulated as:
$Y=Sigmoid(A_{2}\cdot G(GMP(X),A)),$ (6)
where $G$ denotes the graph attention operation.
Upon inputting the $k$ pairs of object feature maps $(f_{q}^{obj},(f_{r}^{obj})^{i})_{i\in[1,k]}$ into DCGA, M object descriptor vectors and N object descriptor vectors were obtained and denoted as $(d_{q}^{M},(d_{r}^{N})^{i})_{i\in[1,k]}$.
III-E2 Object Matching Module
Motivation: In the object matching problem of logical anomaly detection, the correspondences between objects in the query image and those in the reference image must adhere to certain physical constraints: i) An object in the query image may have at most one matching counterpart in the reference image; ii) An effective matching model should suppress the matching of objects that are extraneous or should not be present in the query image.
Task Formulation For each pair of object descriptor vectors $(d_{i}^{M},d_{j}^{N})_{i\in[1,M],j\in[1,N]}$, constraints i) and ii) imply that correspondences come from a partial assignment between the two sets of objects, that is, each possible correspondence should have a confidence value. Therefore, we have defined a partial soft assignment matrix $\mathbf{P}\in[0,1]^{M\times N}$ as:
$\mathbf{P}\mathbf{1}_{N}\leq\mathbf{1}_{M}\quad\text{ and }\quad\mathbf{P}^{\top}\mathbf{1}_{M}\leq\mathbf{1}_{N}.$ (7)
Our goal is to devise an object matching module capable of predicting the registration $\mathbf{P}$ from two sets of descriptor features.
Optimal Matching An optimal transport layer [25] is used to extract the object correspondences between $[1,M]$ and $[1,N]$. Specifically, we first compute a score matrix $\mathbf{S}_{i,j}\in\mathbb{R}^{M\times N}$:
$\mathbf{S}_{i,j}=\left\langle d_{i}^{M},d_{j}^{N}\right\rangle,\forall(i,j)\in M\times N,$ (8)
[…]
Through the assignment matrix $\mathbf{P}$, we obtain the detailed matching results for the object descriptor vectors $d_{i}^{M}$, which include the index of each successfully matched descriptor vector and the corresponding index of the reference object descriptor vector to which it was matched. Through the trash bin, we acquire the indices of the query object descriptor vectors that were not matched, along with the index of the closest matching reference object descriptor vector ( In the $\mathbf{P}$ matrix, the index of the maximum value in the row corresponding to that object descriptor vector).
Figure 6: Architecture of the Anomaly Measurement Module (AMM).
III-F Anomaly Measurement Module
For $k$ pairs of object descriptor vectors $(d_{q}^{M},(d_{r}^{N})^{i})_{i\in[1,k]}$, each pair is feeding into Object Matching Module, yielding $k$ assignment matrices $\mathbf{P}$ and corresponding trash bins. Based on all the $\mathbf{P}$ and trash bins, we label the object descriptor vectors $d_{i}^{M},i\in[1,M]$ of the query image as either matched $d_{i}^{Matched},i\in[1,Matched]$ or unmatched $d_{i}^{Unmatched},i\in[1,Unmatched]$ ($Matched+Unmatched=M$).
To enhance the robustness of the detection results produced by the proposed SAM-LAD, we do not simply use the object masks corresponding to the $d_{i}^{Unmatched},i\in[1,Unmatched]$ as the final score map. Instead, we build an Anomaly Measurement Module (AMM) to detect anomalies for each object in the query image individually, resulting in an overall anomaly score map.
Specifically, based on the indices from the $\mathbf{P}$ and trash bins, we calculate the difference of feature distribution for the object feature maps that were successfully matched with the corresponding object feature maps in the $k$ reference images to create the matching score map. We also calculate the difference of feature distribution for the unmatched object feature maps with the most closely matched object feature maps in the $k$ reference images to create the non-matching score map. The final score map was then obtained by adding these two score maps together. The advantage of this approach is that after matching at the object level, subsequent pixel-level computations allow not only the detection of logical anomalies but also a more detailed check for structural anomalies within each class of object.
A statistical-based estimator is built to estimate the normal distribution of the $k$ feature maps that an object of the query image is matched with, which uses multivariate Gaussian distributions to get a probabilistic representation of the normal class. Suppose a feature map is divided into a grid of $(x,y)\in[1,H]\times[1,W]$ positions where $H\times W$ is the resolution of the feature map used to estimate the normal distribution. At each patch position $(x,y)$, let $F_{xy}=\{f_{xy}^{i},i\in[1,k]\}$ be the normal features from $k$ reference object feature maps. By the assumption that $F_{xy}$ is generated by $(\mu_{xy,\Sigma_{xy}})$, that sample covariance is:
$\Sigma_{xy}=\frac{1}{k-1}\sum_{i=1}^{k}\left(f_{xy}^{i}-\mu_{xy}\right)\left(f_{xy}^{i}-\mu_{xy}\right)^{\mathrm{T}}+\epsilon I,$ (9)
where $\mu_{xy}$ is the sample mean of $F_{xy}$, and the regularization term $\epsilon I$ makes the sample covariance matrix full rank and invertible. Finally, each possible patch position is associated with a multivariate Gaussian distribution. During inference, a query object feature map that is out of the normal distribution is considered an anomaly. For a query object feature map, we use the Mahalanobis distance $\mathcal{M}(f_{xy})$ to give an anomaly score to the patch in position $(x,y)$, where
$\mathcal{M}\left(f_{xy}\right)=\sqrt{\left(f_{xy}-\mu_{xy}\right)^{T}\Sigma_{xy}^{-1}\left(f_{xy}-\mu_{xy}\right)}.$ (10)
The matrix of Mahalanobis distances $\mathcal{M}=\left(\mathcal{M}\left(f_{xy}\right)\right)_{1\leqslant x\leqslant H,1\leqslant y\leqslant W}$ forms an anomaly map.
For object of the $d_{i}^{Matched},i\in[1,Matched]$, matching score map is:
$\mathcal{M}_{Matching}=\sum_{i=1}^{Matched}\mathcal{M}_{i}.$ (11)
For object of the $d_{i}^{Unmatched},i\in[1,Unmatched]$, non-matching score map is:
$\mathcal{M}_{Non-matching}=\sum_{i=1}^{Unmatched}\mathcal{M}_{i}.$ (12)
The final anomaly score $\mathcal{M}_{final}$ of the entire query image is:
$\mathcal{M}_{final}=\mathcal{M}_{Matching}+\mathcal{M}_{Non-matching}$ (13)
In a nutshell, the process of the AMM is summarized in Fig. 6, which is an example of the matched object and the unmatched object. For the remaining object, apply the same forward process as depicted in the example of Fig. 6.
Figure 7: Datasets used in the experiments. (a) MVTec LOCO AD [11]. (b) MVTec AD [1]. (c) DigitAnatomy dataset [35]. Among these datasets, the majority of the MVTec AD consists of structural anomalies, while the MVTec LOCO AD dataset includes both logical and structural anomalies, and the DigitAnatomy dataset includes logical anomalies. The anomalies are annotated by red circles.
IV Experimental verification
In this section, we will conduct comprehensive experiments to validate the effectiveness of the proposed SAM-LAD. Specifically, we will compare it with existing methods on benchmarks from multiple scenarios and perform further analysis to verify the reasons behind the performance of the framework. Finally, we will conduct ablation studies to further analyze the framework’s performance.
IV-A Datasets
In our experiments, we primarily use three public unsupervised anomaly detection datasets and a selection of representative samples from these datasets is illustrated in Fig. 7.
TABLE I: Comprehensive Comparison Results of the Proposed SAM-LAD and Existing Methods. Training indicates whether a method requires retraining with the dataset of a new scene when detecting anomalies in that scene. ()($\uparrow$)( ↑ ) indicates that higher values represent better performance and ()($\downarrow$)( ↓ ) indicates that lower values represent better performance. Bold font indicates the best results, while underlined represents the second-best results.
Methods Indicators Configure Capability Efficency MVTec LOCO AD MVTec AD
[…]
AE[36] $\checkmark$ CNN $\checkmark$ $\times$ 5.0 251.1 57.4 37.8 71.0 80.4
f-AnoGAN[37] $\checkmark$ CNN $\checkmark$ $\times$ 7.7 133.4 64.3 33.4 65.8 76.2
SPADE[38] $\checkmark$ CNN $\checkmark$ $\times$ - 0.9 74.0 45.1 85.5 96.5
Padim[39] $\checkmark$ CNN $\checkmark$ $\times$ - 4.6 78.0 52.1 95.5 96.7
Patchcore[10] $\checkmark$ CNN $\checkmark$ $\times$ 11.4 25.1 83.5 34.3 99.1 98.1
THFR[14] $\checkmark$ CNN $\checkmark$ $\checkmark$ - 7.69 86.0 74.1 99.2 98.2
DSKD[15] $\checkmark$ CNN $\checkmark$ $\checkmark$ - - 84.0 73.0 - -
GLCF[13] $\checkmark$ Transformer $\checkmark$ $\checkmark$ 52.6 11.2 83.1 70.3 98.6 98.2
SAM-LAD(ours) $\times$ Transformer $\checkmark$ $\checkmark$ 34.6 1.9 90.7 83.2 98.0 97.7
MVTec LOCO AD: The MVTec LOCO AD dataset [11] was recently released by MVTec Software GmbH, which is specifically developed for logical anomalies and comprises five object categories, each containing both structural and logical anomalies in the test set. It has a total of 2,076 anomaly-free samples and 1,568 samples for testing. Each of the 1,568 test images is either anomaly-free or contains at least one structural or logical anomaly. Specifically, logical anomalies primarily include missing elements, extra items, disordered sequences, and inappropriate combinations, among others. Pixel-level annotations are provided as the ground truth for testing.
MVTec AD: The MVTec AD [1] comprises 10 object categories and 5 texture categories, with a total of 4,096 anomaly-free samples and 1,258 anomaly samples in the testing set. The anomaly types include only local structural damage. Pixel-level annotations are provided as the ground truth for testing.
DigitAnatomy: In a recent study[35], a groundbreaking synthetic logical dataset was introduced, consisting of digits organized in a grid pattern. Images containing digits in the correct sequential order were deemed normal, while those with deviations were categorized as abnormal. The dataset comprises a variety of simulated anomalies, including missing digits, out-of-sequence digits, flipped digits, and zero digits. These types of anomalies exhibit a greater degree of logical patterns. Image-level annotations are provided as the ground truth for testing.
IV-B Evaluation metrics
We use the Area Under the Receiver Operating Characteristic Curve (AUROC) score as a threshold-free metric to evaluate image-level anomaly detection. For anomaly localization in both MVTec AD and MVTec LOCO AD datasets, AUROC is also suitable for assessing structural anomalies. However, logical anomalies of the MVTec LOCO AD, e.g. a missing object, are challenging to annotate and segment on a per-pixel basis. To evaluate anomaly localization performance, we use the saturated Per-Region Overlap (sPRO) metric[11] with the per-pixel false-positive rate of 5%, which is a generalized version of the PRO metric [1]. This metric reaches saturation once the overlap with the ground truth achieves a predefined saturation threshold. All thresholds are also provided by MVTec LOCO AD dataset.
IV-C Implementation Details
In our experiments, we resize each image to a resolution of 224 $\times$ 224 and normalize the pixel intensities based on the mean value and standard deviation obtained from the ImageNet dataset[40]. Additionally, the $k$-nearest is set to 2. All experiments were conducted on a computer equipped with Xeon(R) Gold 6230R CPUs@2.60GHZ and one NVIDIA A100 GPU with 40GB of memory.
IV-D Comparison With the State-of-the-Art Models
In this subsection, the proposed SAM-LAD framework is analyzed in comparison with several state-of-the-art (SoTA) methods on several benchmark datasets.
IV-D1 Comprehensive Comparison
In the initial phase of our study, a comprehensive evaluation was conducted to compare the proposed SAM-LAD with existing methods, including AE[36], f-AnoGan[37], SPADE[38], Padim[39], Patchcore[10], THFR[14], DSKD[15], and GLCF[13]. The results are presented in Table I. Notably, all existing methods require additional training for different scenarios, resulting in limited generalization. In contrast, our method achieves zero-shot capabilities, allowing for plug-and-play functionality in any scene while ensuring optimal logical detection performance. Analyzing the network structures employed in these methods, it was observed that most existing approaches are built upon convolutional neural networks (CNNs), with only a few utilizing vision transformers, which are still in the early stages of exploration. A significant limitation of CNN-based models lies in their inability to capture global semantics effectively, which is crucial for logical anomaly detection.
In terms of inference efficiency and computational requirements, our SAM-LAD employs a transformer structure and upsampled feature maps, which affects the inference time compared to other schemes. However, the FLOPs of SAM-LAD is still within acceptable limits. When considering benchmark performance, the SAM-LAD demonstrates advanced capabilities in Mvtec AD for detecting structural anomalies and Mvtec LOCO AD for identifying logical anomalies, owing to its proficiency in robust segment and object matching.
TABLE II: Quantitative detection and localization results of the SAM-LAD framework on the MVTec LOCO AD dataset. Results for each category are given as logical anomalies/structural anomalies or the average of both. Overall averages are given as logical anomalies/structural anomalies and the average of both. The results of the comparison methods are from [11], [13], and [41].
[…]
AE
[36]
VAE
[42]
f-AnoGAN
[37]
MNAD
[43]
EfficientAD-S
[29]
GCAD
[11]
THFR
[14]
DSKD
[15]
GLCF
[13]
Image-Level AUROC Breakfast box 58.0/47.7 47.3/38.3 69.4/50.7 59.9/60.2 - 87.0/80.9 78.0 - 86.7/79.1 96.7/85.2
Juice bottle 67.9/62.6 61.3/57.3 82.4/77.8 70.5/84.1 - 100/98.9 97.1 - 98.7/93.3 98.7/96.5
Pushpins 62.0/66.4 54.3/75.1 59.1/74.9 51.7/76.7 - 97.5/74.9 73.7 - 80.1/78.6 97.2/79.2
Screw bag 46.8/41.5 47.0/49.0 60.8/56.8 46.8/59.8 - 56.0/70.5 88.3 - 80.1/78.6 95.2/77.9
Splicing Connectors 56.2/64.8 59.4/54.6 68.8/63.8 57.6/73.2 - 89.7/78.3 92.7 - 89.6/89.7 91.4/88.6
Average 58.2/56.6 57.4 53.8/54.8 54.3 65.9/62.7 64.3 60.1/70.2 65.1 94.1/85.8 90.0 86.0/80.7 83.4 86.0 81.2/86.9 84.0 82.4/83.8 83.1 95.8/85.5 90.7
Pixel-Level sPRO Breakfast box 18.9 16.5 22.3 8.0 - 50.2 58.3 56.8 52.8 81.9/79.1
Juice bottle 60.5 63.6 56.9 47.2 - 91.0 89.6 86.5 91.3 94.4/93.5
Pushpins 32.7 31.1 33.6 35.7 - 73.9 76.3 82.5 61.5 76.2/74.2
Screw bag 28.9 30.2 34.8 34.4 - 55.8 61.5 62.7 61.5 86.3/71.6
Splicing Connectors 47.9 49.6 19.5 44.2 - 79.8 84.8 76.7 78.5 89.1/85.2
Average 46.0/29.6 37.8 45.9/30.5 38.2 46.0/20.9 33.4 26.6/41.2 33.9 74.8/ 80.8 77.8 71.1/69.2 70.1 74.1 73.0 70.0/70.6 70.3 85.6/ 80.7 83.2
  • 1
    The best performance is indicated by bold font, while the second best is indicated by an underline.
Figure 8: Examples of qualitative detection results for logical and structural anomalies using our SAM-LAD framework on the MVTec LOCO AD dataset. The Ground truth defects are annotated by red circles.
TABLE III: The AUROC Results of Various Methods in MVTec AD at the Image/Pixel-level
[…]
AE
[36]
PMB-AE
[9]
MKD
[44]
RIAD
[45]
DRAEM
[3]
RD4AD
[46]
Padim
[39]
Patchcore
[10]
C-FLOW
[47]
GLCF
[13]
SAM-LAD (ours)
Carpet 67.0/87.0 93.1/92.3 79.3/95.6 84.2/94.2 97.0/95.5 98.9/98.9 99.8/98.9 98.7/99.0 99.8/98.9 99.8/98.2 100/99.3
Grid 69.0/94.0 97.1/94.3 78.1/91.8 93.0/85.8 99.1/96.8 99.2/95.3 99.2/93.6 99.2/95.0 99.1/96.7 99.3/94.8 99.2/98.6
Leather 46.0/78.0 94.5/96.7 95.1/98.1 100/99.4 100/98.6 100/99.4 100/99.1 100/99.3 100/99.1 100/99.0 100/98.9
Tile 52.0/59.0 97.2/90.7 91.6/82.8 98.7/89.1 99.6/99.2 99.3/95.6 98.1/91.2 98.7/95.6 98.1/91.2 99.8/95.1 100/96.3
Wood 83.0/73.0 100/86.5 94.3/84.8 99.6/96.3 99.9/99.3 100/99.3 96.7/94.9 98.2/98.7 96.7/94.9 99.7/98.9 97.3/94.3
Bottle 88.0/93.0 93.7/95.2 99.4/96.3 99.9/98.4 99.2/99.1 100/98.7 99.9/98.1 100/98.6 100/99.0 100/98.4 100/98.6
Cable 61.0/82.0 85.6/94.2 89.2/82.4 81.9/84.2 91.8/94.7 95.0/97.4 92.7/95.8 99.5/98.4 97.6/97.6 100/98.2 95.2/97.5
Capsule 61.0/94.0 82.3./92.1 80.5/95.9 88.4/92.8 98.5/94.3 96.3/98.7 91.3/98.3 98.1/98.8 97.7/99.0 95.5/98.9 96.8/98.2
Hazelnut 54.0/97.0 99.4/92.5 98.4/94.6 83.3/96.1 100/92.9 99.9/98.9 92.0/97.7 100/98.7 100/98.9 100/98.9 99.3/99.4
Meta nut 54.0/89.0 85.8/84.5 82.7/86.4 88.5/92.5 98.7/96.3 100/97.3 98.7/96.7 100/98.4 99.3/98.6 100/97.8 99.9/98.0
Pill 60.0/91.0 86.1/91.1 82.7/89.6 84.5/98.8 93.9/97.6 97.0/99.6 85.8/97.4 98.1/99.4 91.9/98.9 95.3/99.4 98.6/98.2
Screw 51.0/96.0 97.0/97.7 83.3/96.0 100/98.9 100/98.1 99.5/99.1 96.1/98.7 100/98.7 99.7/98.9 92.5/98.8 95.7/96.8
Toothbruth 74.0/92.0 95.8/97.5 92.2/96.1 83.8/95.7 98.9/97.6 96.6/98.2 93.3/94.7 96.6/97.4 96.8/98.9 96.3/98.1 96.3/99.0
Transistor 52.0/90.0 80.8/92.4 85.6/76.5 90.9/87.7 93.1/90.9 96.7/92.5 97.4/97.2 100/96.3 95.2/98.0 100/97.5 95.8/98.5
Zipper 80.0/88.0 77.3/95.4 93.2/93.9 98.1/97.8 100/98.8 98.5/98.2 90.3/98.2 99.4/98.5 98.5/99.1 97.2/97.7 96.2/94.1
Average 63.0/87.0 91.8/92.1 87.7/90.7 91.7/94.2 98.0/97.3 98.5/97.8 95.5/96.7 99.1/98.1 98.3/98.6 98.3/98.0 98.0/97.7
IV-D2 MVTec LOCO AD
We compare our proposed SAM-LAD framework with existing methods, including baseline approaches such as f-AnoGAN[37], AE[36], VAE[42], and MNAD[43], as well as the top 5 best-performing SoTA methods on MVTec LOCO AD dataset’s leaderboard on ”Papers with Code [^1] https://paperswithcode.com/sota/anomaly-detection-on-mvtec-loco-ad”: EfficientAD-S[29], GCAD[11], THFR[14], DSKD[15], and GLCF[13].
The comparative results are presented in Table II. Compared to the best baseline method and the SoTA method, our framework improves +25.6 and +0.7 on image-level AUROC, +45.0 and +5.4 on pixel-level sPRO, respectively. It is worth mentioning that, among all the comparison methods, only our approach implemented zero-shot while simultaneously maintaining the best detection performance. Furthermore, Fig. 8 displays several detection results of SAM-LAD on both structural and logical anomalies within the MVTec LOCO AD dataset, illustrating that our framework is adept at precisely pinpointing both types of anomalies.
Fig. 10 depicts the efficacy of each method (EfficientAD-M, GCAD, SPADE[38], DRAEM[3], GLCF) on the MVTec LOCO AD dataset concerning both structural and logical anomalies. The results indicate that our proposed strategy showcases strong capabilities in identifying both structural and logical anomalies. It is noteworthy to mention that, in comparison to the current best SoTA method, the pixel-level detection performance for logical anomalies has witnessed a significant increment of 9.1 through our framework. This highlights our method’s superior proficiency in discerning logical anomalies, which aligns perfectly with our foundational intent. It enables an in-depth comprehension of the logical interrelations among each object within the entire scene and proficiently identifies the corresponding logical discrepancies.
IV-D3 MVTec AD
Beyond affirming the performance of our SAM-LAD in detecting logical anomalies, we have further appraise its capability to identify structural anomalies on the MVTec AD dataset, which exclusively encompasses structural anomalies. To be specific, given that each subset within the MVTec AD dataset features a scene with a single object against a regular background, we facilitate the segmentation into this single object by employing the SAM filter for subsequent matching and computation. Owing to the one-to-one fixed matching paradigm, the ultimate detection results are directly yielded by the AMM, thereby serving as an indirect assessment of the efficacy of our proposed AMM.
We compare SAM-LAD with several SoTA methods, which are AE-SSIM[36], PMB-AE[9], MKD[44], RD4AD[46], RIAD[45], DRAEM[3], Padim[39], Patchcore[10], C-FLOW AD[47], and GLCF[13]. Table III shows the quantitative comparison results. The proposed SAM-LAD achieves remarkable image-level anomaly detection results and pixel-level anomaly localization results, obtaining an AUROC of 98.0/97.7 across 15 categories. Remarkably, SAM-LAD exhibits significantly better performance compared to baseline methods. Moreover, when compared with the SoTA methods such as Patchcore and C-FLOW, our framework achieves comparable detection and localization accuracy. A selection of qualitative results on the MVTec AD dataset is depicted in Fig. 9.
In conclusion, the proposed SAM-LAD showcases exceptional performance in the context of industrial anomaly detection. Fig. 11 offers a comprehensive comparison between SAM-LAD and existing methods regarding their image-level and pixel-level detection capabilities on the MVTec AD and MVTec LOCO AD datasets. The results unequivocally indicate that our framework attains the highest precision in anomaly detection and localization across these benchmarks, thereby evidencing the efficacy and versatility of our SAM-LAD in addressing an array of anomalies, including localized structural anomalies and complex logical anomalies.
Figure 9: Examples of qualitative detection results for structural anomalies using our SAM-LAD framework on the MVTec AD dataset.
IV-D4 DigitAnatomy
To further validate the efficacy of the proposed SAM-ALD in detecting logical anomalies, we conduct a comparative experiment employing the DigitAnatomy dataset. A gamut of comparative assessments was performed, utilizing an array of methods such as AE[36], GANomaly[48], f-AnoGAN[37], and SQUID[35], coupled with the integrative Fastflow[49] and Patchcore[10], in conjunction with Draem[3], MKD[44], RD4AD[46], and GLCF[13].
The comparative results are delineated in Table IV and the qualitative results are shown in Fig. 12. The results show that our framework markedly transcends existing methods. In particular most efficacious GLCF method, our novel SAM-LAD method realizes a substantial enhancement, accruing a gain of +13.7 in AUROC. Albeit the SQUID method was devised alongside the DigitAnatomy dataset, it garners merely an unassuming AUROC of 55.7. Conversely, our SAM-LAD framework attains a formidable AUROC of 92.3, thereby accentuating its preeminent efficacy.
IV-E Further Analysis
The outstanding logical anomaly detection results of our framework are based on the good performance of the Object Matching Model (OMM). Therefore, to verify the reasons behind the exceptional performance, for each category in MVTec LOCO AD and DigitAnatomy dataset, we compile a confusion matrix of the object matching results in Table V and Table VI, respectively. The ground truth is categorized by the number of objects in each test image that should be correctly matched and those that should not be matched. We use precision, recall, and F1 score as the metrics to evaluate the OMM’s performance.
The results in Table V and Table VI demonstrate the strong matching capability of the OMM, validating its effectiveness. However, in Table V, the matching accuracy for the pushpins and splicing connectors categories drops significantly. This is because, in these scenarios, each object is very similar, posing a considerable challenge for the object matching model. However, we did not directly use the unmatched object masks as the anomaly score. Instead, we design the AMM (Anomaly Measurement Module) to further calculate anomalies in the objects. This novel design greatly reduces the proposed SAM-LAD framework’s dependency on matching accuracy, as demonstrated by the final anomaly detection results in Section IV-D. A selection of the representative samples from the visualized matching results is shown in Fig. 13.
TABLE IV: Quantitative AUROC Comparison Results on the DigitAnatomy Dataset
[…]
AE
[36]
GANomaly
[48]
f-AnoGAN
[37]
SQUID
[35]
Fastflow
[49]
Patchcore
[10]
DRAEM
[3]
MKD
[44]
RD4AD
[46]
GLCF
[13]
SAM-LAD(ours)
AUROC 50.2 62.9 54.3 55.7 56.2 59.4 52.6 54.8 58.5 78.6 92.3
IV-F Ablation Experiment
IV-F1 Impact of the Backbone
We use the pre-trained backbone as the feature extractor in the proposed SAM-LAD. In this study, we primarily considered three different pre-trained backbone networks: the Wide-ResNet50 of the CNN type, the Swin-transformer of the ViT type, and the DINOv2 of the ViT type. The main results are presented in Table VII, where it is distinctly that DINOv2 manifests the most superior performance. Therefore, we adopt DINOv2 as the feature extractor for the proposed framework.
IV-F2 Impact of the FeatUp Configuration
Upon obtaining the feature maps from the images, we implement the FeatUp operation to upsample these maps and thus restore the spatial information that had been lost. In the study of FeatUp[18], the authors proffered four configurations of upsampling: 2$\times$, 4$\times$, 8$\times$, 16$\times$, and 32$\times$, which are displayed in Fig. 14. To verify the efficacy of FeatUp and to assess the impact of different upsampling configurations on the performance of anomaly detection, we conduct an ablation study focusing on the upsampling factors.
Figure 10: Comparison results of structural anomalies and logical anomalies on MVTec LOCO AD dataset.
Figure 11: Image and pixel-level performance comparison of the proposed SAM-LAD and existing methods on two benchmarks
Figure 12: Example of qualitative detection results for logical anomalies using our SAM-LAD framework on the Digitanatomy dataset. The anomalies are annotated by red circles.
Figure 13: The matching results for each subset in MVTec LOCO AD and DigitAnatomy dataset, where the green lines indicate matching objects.
The main results are illustrated in the Table VIII. In comparison with the original feature maps, FeatUp significantly enhances detection capabilities. More specifically, as the factor of upsampling escalates, the detection and segmentation performance of SAM-LAD are concurrently fortified. Nevertheless, when the feature maps are subjected to upsampling by 16$\times$ and 32$\times$, there is dramatic inflation in the model’s FLOPs and a precipitous deterioration in inference speed, circumstances that are indefensible in pragmatic industrial settings. Consequently, to harmonize detection performance with the imperative of real-time inference, an 8$\times$ upsampler was opted for in our framework to restore the spatial information lost in the feature maps for use in subsequent computational processes.
IV-F3 Impact of the DCGA
To validate the effectiveness of the proposed DCGA, an ablation experiment was conducted without the incorporation of DCGA. The inception of DCGA was intended to extract high-dimensional features into a single feature vector, thereby simplifying the computation of similarity between objects. Pooling represents the most extensively utilized method for feature extraction. Consequently, we compare DCGA with both global average pooling (GAP) and global max pooling (GMP), with results delineated in Table IX. It can be observed that GMP outperforms GAP because only certain regions of the object feature map possess discernible feature values, while the background is predominantly nullified. Therefore, in such scenarios, GMP is more effective in capturing object features. Nevertheless, simply selecting the maximum value for pooling is insufficient. The introduction of DCGA has greatly enriched the description of object features. Utilizing the principles of graph neural networks, DCGA not only extracts efficacious object features from the high-dimensional channels but also captures the interrelationship information of features among objects within the same scene. This substantially improves the matching capabilities of the optimal-matching model and the ultimate performance of anomaly detection, thus confirming the validity of the proposed DCGA.
TABLE V: The Precision, Recall, and F1 Score of the Object Matching Module in the MVTec LOCO AD dataset.
breakfast box
Actual Predicted Positive Negative Precision: 99.2%
Positive 1502 19 Recall: 98.8%
Negative 12 117 F1 Score: 99.0%
juice bottle
Actual Predicted Positive Negative Precision: 99.5%
Positive 1503 26 Recall: 98.3%
Negative 8 110 F1 Score: 98.9%
pushpins
Actual Predicted Positive Negative Precision: 89.0%
Positive 3336 505 Recall: 86.9%
Negative 413 510 F1 Score: 87.9%
screw bag
Actual Predicted Positive Negative Precision: 99.5%
Positive 2426 28 Recall: 98.9%
Negative 13 261 F1 Score: 99.2%
splicing connectors
Actual Predicted Positive Negative Precision: 91.1%
Positive 2471 339 Recall: 87.9%
Negative 242 508 F1 Score: 89.5%
Mean Precision 95.66%
Mean Recall 94.16%
Mean F1 Score 96.52%
TABLE VI: The Precision, Recall, and F1 Score of the Object Matching Module in the DigitAnatomy.
Actual Predicted Positive Negative Precision: 96.5%
Positive 2853 147 Recall: 95.1%
Negative 102 1398 F1 Score: 95.8%
TABLE VII: Ablation Experiments of the Different Backbone on the MVTec LOCO AD Dataset.
Metrics Network Wide-ResNet50 Swin-Transformer DINOv2
Det.(AUC) 82.3 85.8 90.7
Seg.(sPRO) 72.9 79.7 83.2
Figure 14: Visualizing PCA components with different upsampling configurations of the FeatUp.
TABLE VIII: Ablation experiments of the Different FeatUp’s Upsampling Factors on the MVTec LOCO AD Dataset.
[…]
Performance Det.(AUC) 81.4 83.1 86.2 90.7 91.2 91.8
Seg.(sPRO) 71.1 74.1 79.7 83.2 83.8 84.3
Efficency FLOPs(Gb) 9.2 15.2 23.8 34.6 65.1 94.2
FPS 7.2 5.5 3.4 1.9 0.9 0.1
TABLE IX: Ablation Experiments of the Different Feature Compress Methods on the MVTec LOCO AD Dataset.
[…]
Object Matching Precision 90.65% 92.26% 95.66%
Recall 89.02% 91.43% 94.16%
F1 Score 89.83% 91.84% 96.52%
Anomaly Detecting Det.(AUC) 85.3 86.2 90.7
Seg.(sPRO) 74.7 76.2 83.2
V Conclusion
In this paper, we novelty propose a zero-shot framework called SAM-LAD to address logical anomaly detection in complex scenes. We introduce a pre-trained SAM to obtain masks for all objects in the query image. Utilizing the pre-trained DINOv2 and FeatUp operations, we derive the upsampled feature map. Through nearest neighbor search on the query image, we obtain a reference image and its corresponding upsampled feature map. By sequentially combining object masks with upsampled feature maps, we acquire each object feature maps in both the query and reference images. Subsequently, we regard each object as a key point and employed the proposed DCGA mechanism to efficiently compress each object’s feature map into a feature vector. Then, we propose the OMM, matching all object feature vectors from the query image with those in the reference image to obtain a matching matrix. Finally, based on the matching matrix, we propose AMM to compute the distribution difference estimation of the matched objects’ features, resulting in the final anomaly score map.
Nevertheless, within the scope of our investigation, a limitation has come to light, that could potentially impede the overall performance of our framework. Specifically, when objects in the scene all belong to the same category, due to their high similarity, the object matching model exhibit some false positives, such as the pushpins shown in Table V. Consequently, as part of our future work, we aim to explore potential enhancements to mitigate these issues and further boost the performance of our framework.